A 2500 kg delivery truck contains 125 cavans of rice, each of mass 50 kg. It is moving at a speed of 5.0 m/s. How many cavans of rice must be unloaded at a nearby grocery so that the ruck can move at 9.0 m/s and yet maintain the same momentum?


1
Expert's answer
2021-01-12T12:12:07-0500

Let's first find the initial momentum of the delivery truck:


"p_i=(M+mN)v,""p_i=(2500\ kg+50\ kg\cdot 125)\cdot 5.0\ \dfrac{m}{s}=43750\ \dfrac{kgm}{s}."

As we know from the conditions of the question, "p_i=p_f=p", so we can write:


"p=(M+mN)v,""N=\dfrac{p-Mv}{mv},""N=\dfrac{43750\ \dfrac{kgm}{s}-2500\ kg\cdot 9\ \dfrac{m}{s}}{50\ kg\cdot 9\ \dfrac{m}{s}}=47\ cavans."


Therefore, "N_{unload}=125-47=78\ cavans".

Answer:

"N_{unload}=78\ cavans".